OpenStudy (anonymous):

Related Rates problem... A puddle is evaporating in such a way that its diameter is decreasing at a rate of 0.1 cm/min. At what rate is the area of the puddle decreasing when the diameter is 8cm.

OpenStudy (anonymous):

ooohhh funn =]

OpenStudy (anonymous):

noo.. I'm horrible at these and word problems in general.

OpenStudy (abb0t):

The area of a circle is \[A = \Pi r^2\]

OpenStudy (anonymous):

relevant equation \[A=\frac{\pi}{4}d^2\]

OpenStudy (anonymous):

you're trying to fnd \[\frac{dA}{dt}=?\] you know the rate of change of thediameter \[\frac{dD}{dt}\] lol double d=]

OpenStudy (anonymous):

if you take the derivative of the original equation.. .what do you get =]

OpenStudy (abb0t):

Your answer should be: \[-4\Pi\] try it first and let me know how it goes.

OpenStudy (anonymous):

\pi is the code for pi =]...

OpenStudy (anonymous):

ahh, hold on ;/

OpenStudy (anonymous):

\[ \pi \ ] without the bracket being out in the middle of nowhere

OpenStudy (abb0t):

It might help if you write down all your givens first, monroe. That will help guide you in the right direction.

OpenStudy (anonymous):

A'=(pi/4)*2d ?

OpenStudy (abb0t):

You already know your dr/dt

OpenStudy (anonymous):

almost when you took the derivative of \[d^2\] in terms of time... you forgot something... there was no t.... soooooo

OpenStudy (anonymous):

in terms of time>?

OpenStudy (abb0t):

ok...

OpenStudy (anonymous):

yes \[\frac{d}{dt}\]if you had something like this \[\frac{d}{dt}[t^2]=2t\]. however you have \[\frac{d}{dt}(d^2)\] we can't just say it's 2d ... think about when you implicitly differentiated before

OpenStudy (anonymous):

so because it's d/dt the d changes to t?

OpenStudy (anonymous):

ughhh.. okay soo.. 2d(d/dt)

OpenStudy (anonymous):

\[\frac{dD}{dt}\]...

OpenStudy (anonymous):

ughhh.. ;[

OpenStudy (anonymous):

A'=\[\frac{ \pi }{ 4 }*2d \frac{ dD }{ dt }\]

OpenStudy (anonymous):

I didn't see that you wrote "you know the rate of change of thediameter dD/dt" at the top.

OpenStudy (anonymous):

think back when you did chain rule what we are noting is that the d in our equation is some sort of function of time otherwise \[d=f(t)\] so going back to chain rule and the way i showed you earlier if \[d=f(t)\] and \[y=cd^2 \] using chain rule, \[\frac{dy}{dt}=\frac{dd}{dt}*\frac{dy}{dd}\]

OpenStudy (anonymous):

so whenever you differentiate in respects to a term that is not within the function itself you have to think of all other terms as a function of time

OpenStudy (anonymous):

and yep what you said was correct.. and now you have everything you need to solve ... just plug in and you'll get the rate at which your area is changing =] fun huh?

OpenStudy (anonymous):

what did you get

OpenStudy (abb0t):

I have no idea what he's talking abt. Lol

OpenStudy (anonymous):

A'=\[\frac{ \pi }{ 4 }2*8*\] wait is 0.1 dD/dt?

OpenStudy (abb0t):

\[\frac{ dA }{ dt } = 2\pi4(0.1)\]

OpenStudy (anonymous):

-.1

OpenStudy (anonymous):

yeahh yea.. because it's decreasing right>?

OpenStudy (anonymous):

yes dD/dt = the rate of change of your Diameter in respect to time or \[\frac{\Delta D}{\Delta t}\]

OpenStudy (anonymous):

-0.4pi cm^2/min :)

OpenStudy (anonymous):

and btw i do not know how you get 4 for 2d thats why you don't do it with radius you have to multiply by 2 to get it back to diameter =]

OpenStudy (anonymous):

... me?

OpenStudy (anonymous):

oh mkay.

OpenStudy (abb0t):

lol

OpenStudy (anonymous):

actually i take that back you can't even do that... the rate of radius is much different than the rate of diameter